1. Apple stock is selling for $120 per share. Call options with a $117 exercise price are priced at $12. What is the intrinsic value of the option, and what is the time value?  

2. Twitter is trading at $34.50. Call options with a strike price of $35 are priced at $2.30. What is the intrinsic value of the option, and what is the time value?  
Answer #1

1

Intrinsic value = selling price-exercise price = 120-117=3

Time value = option price-intrinsic value = 12-3 = 9

2

Intrinsic value = Max(selling price-exercise price,0) = Max(34.5-35,0) = Max(-0.5,0) = 0

Time value = option price-intrinsic value = 2.3-0 = 2.3

3